# Hierarchical Importance Sampling for Estimating Occupation Time for SDE Solutions

### v1

@@ Abstract
This study considers the estimation of the complementary cumulative distribution function of the occupation time (i.e., the time spent below a threshold) for a process governed by a stochastic differential equation. The focus is on the right tail, where the underlying event becomes rare, and using variance reduction techniques is essential to obtain computationally efficient estimates. Building on recent developments that relate importance sampling (IS) to stochastic optimal control, this work develops an optimal single-level IS (SLIS) estimator based on the solution of an auxiliary Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE). The cost of solving the HJB-PDE is incorporated into the total computational work, and an optimized trade-off between preprocessing and sampling is proposed to minimize the overall cost. The SLIS approach is extended to the multilevel setting to enhance efficiency, yielding a multilevel IS (MLIS) estimator. A necessary and sufficient condition under which the MLIS method outperforms the SLIS method is established, and a common-likelihood MLIS formulation is introduced that satisfies this condition under appropriate regularity assumptions. The classical multilevel Monte Carlo complexity theory can be extended to accommodate settings where the single-level variance varies with the discretization level. As a special case, the variance-decay behavior observed in the IS framework stems from the zero-variance property of the optimal control. Notably, the total work complexity of MLIS can be better than an order of two. Numerical experiments in the context of fade duration estimation demonstrate the benefits of the proposed approach and validate these theoretical results.
Keywords: Rare-event estimation, occupation time, stochastic differential equation, importance sampling, multilevel Monte Carlo, Hamilton–Jacobi–Bellman, variance decay.
@@ 1 Introduction
This study considers a system in which a process evolves according to a general stochastic differential equation (SDE), with the goal of characterizing the distribution of the occupation time, specifically, the time a process spends below a critical threshold. Occupation time has been analytically studied in the context of Markov chains using matrix-based techniques [1]. In contrast, the present work addresses a more general setting in which the underlying dynamics are governed by SDEs with continuous state spaces, requiring advanced stochastic-simulation techniques for accurate estimation.
@@
Among such techniques, importance sampling (IS) [6] is widely used for rare-event estimation in SDEs. Several studies have established a connection between IS and stochastic optimal control (SOC) for standard SDEs [7, 8, 9, 2, 10, 11], where the optimal change of measure is characterized by the solution of an auxiliary Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE). For example, [9] develops an optimal change of measure via SOC to estimate rare-event probabilities for McKean–Vlasov SDEs, while [2] applies a similar approach to estimate the CCDF of the fade duration in wireless communications. Although these studies have demonstrated substantial variance reduction and have analyzed the sampling cost of IS estimators, they do not account for the computational cost of obtaining the optimal change of measure because they solve the required auxiliary HJB-PDE off-line.
In the present study, the trade-off between the accuracy of the control (10) and the resulting variance reduction is investigated, and the cost of solving the auxiliary HJB-PDE (11) is incorporated into a comprehensive analysis of the total computational work of the IS estimator. The analysis reveals that the variance of the IS estimator decreases with the target accuracy, resulting in a lower overall computational workload for IS compared to the crude MC method.
In order to further improve the efficiency of the single-level IS (SLIS) estimator for rare-event probability estimation, we extend our previous work [2] to the multilevel setting by combining IS with the multilevel MC (MLMC) method introduced by [12]. Several studies have investigated combining IS with the MLMC method. For example, one study [13] applied IS to address the high‑kurtosis phenomenon in the MLMC method for stochastic reaction networks, and another study [14, 15] proposed variance‑reduction schemes based on suboptimal, levelwise constant controls. The most recent development [16] combines SOC-based IS with the MLMC method in a single‑likelihood multilevel IS (MLIS) framework, achieving substantial computational savings. In the context of Bayesian inverse problems, various methods, such as those in [17, 18], exploit coarse-level likelihood evaluations to inform fine-level sampling, demonstrating that reusing the likelihood structure across model hierarchies can be highly effective. Building on these insights and following the method in [16], this work introduces a novel common‑likelihood MLIS approach, highlighting its benefits in reducing the total computational work of MLIS.
@@
The remainder of this paper is structured as follows. Section 2 describes the problem setting. Next, Section 3 details the SLIS method, presenting the construction of the optimal estimator, exploring the influence of the accuracy of the auxiliary HJB-PDE on the variance, and reporting the numerical results. Then, Section 4 reviews the MLMC method and provides the sufficient and necessary conditions for its efficiency. Section 5 introduces the MLIS approach, including the single-level likelihood formulation and common-likelihood strategy. Moreover, Section 6 presents the analysis of the effect of the HJB-PDE accuracy on the efficiency of the MLIS method and the proposed optimal work formulation, providing a numerical example and investigating the work rate when the single-level variance decays.
@@ 2 Problem Setting
where $f(x):=\mathbbm{1}_{\{x<\gamma_{\text{th}}\}}$. The aim is to characterize the CCDF of the occupation time $Z(T)$, that is,
@@
where $g_{w}(x)=\mathbbm{1}_{\{x>w\}}$ for a given threshold $w>0$. This formulation can be compactly described by the following coupled system:
@@ 4.2 Sufficient and Necessary Condition for MLMC Advantages
First,  (35) is equivalent to the following inequality:
and simplifying both sides of (36).
Repeating this backward until any level $\ell<L_{\mathrm{opt}}$ yields
This result implies that, starting from any coarse level $\ell$,
This result implies that, starting with a coarse level $\ell_{0}$,
that is, the MLMC method achieves less sampling work than the SLMC method when initialized at level $\ell_{0}$. ∎
@@
$\sqrt{2V_{\ell_{0}+1}}>\sqrt{V_{\ell_{0}}}+\sqrt{3\,V_{\ell_{0}+1,\ell_{0}}}.$
Dividing this result throughout by $\sqrt{V_{0}}$ yields
By Assumption (1),
Rearranging this result yields (45), completing the proof. ∎
@@ 5.1 Single-Level Likelihood
$\mathcal{A}_{\mathrm{MLIS}}=\frac{1}{M_{0}}\sum_{m=1}^{M_{0}}\tilde{g}_{w}^{(0,m)}+\sum_{\ell=1}^{L}\frac{1}{M_{\ell}}\sum_{m=1}^{M_{\ell}}\left(\tilde{g}_{w}^{\ell}-\tilde{g}_{w}^{\ell-1}\right)^{(m)},$ (48)
where
$\tilde{g}_{w}^{\ell}=g_{w}\left(\tilde{Z}_{N_{\ell}}^{\ell}\right)L_{N_{\ell}}.$ (49)
@@ 5.1.1 Nonsmooth Case
The same application described in Section 3.3 is considered, where the observable $g_{w}$ and drift $f$ are nonsmooth functions. This work investigates the performance of the proposed MLIS estimator (48) by fixing an arbitrary coarse level $N_{0}=20$ and computing the variance $V^{\mathrm{IS}}_{\ell,\ell-1}$ for values of $\ell$ to assess whether the condition in (46) can be satisfied. This analysis is repeated for various levels of the auxiliary HJB-PDE accuracy, following the same approach as earlier to study the behavior of $V_{N}^{\mathrm{IS}}$ (see Section 3.3). Figure 4 presents the results.
@@
Figure 4 reveals that no prospect exists for MLIS to outperform SLIS when $P\geq 80$. The only scenario in which MLIS may offer a slight advantage is when $P=40$. Even then, the advantage is only for tolerances requiring numerous levels, specifically, when $\mathrm{TOL}\leq 0.003$, corresponding to $L_{\mathrm{opt}}\geq 8$.
We compute the sampling work of MLIS using the arbitrary choice $N_{0}=20$ to further investigate this observation. Over the range of relative tolerances $[0.1,0.003]$, SLIS consistently outperforms MLIS for all auxiliary HJB-PDE accuracy values. For a more rigorous comparison, the optimal coarse level $\ell_{0}^{\mathrm{opt}}$ was computed using Algorithm 2, demonstrating that $\ell_{0}^{\mathrm{opt}}=L_{\mathrm{opt}}$ for all considered tolerance values, implying that MLIS reduces to SLIS across the entire tolerance range.
@@ 5.1.2 Smooth Case
This section introduces the smooth approximations $f^{d}$ and $g_{w}^{c}$ for the functions $f$ and $g_{w}$, respectively, to improve the multilevel variance convergence rates. Rather than estimating the original quantity $q_{w}$, the smoothed quantity $q_{w}^{c,d}=\mathbb{E}\left[g_{w}^{c}(Z^{d}(T))\right]$ is approximated, where $Z^{d}(T)=\int_{0}^{T}f^{d}\left(h(\boldsymbol{X}(s))\right)\,ds$. Smoothing introduces an approximation error $\Delta^{c,d}=|q_{w}-q_{w}^{c,d}|$, which can be controlled and is of the order $\sqrt{\mathrm{TOL}}$, where $\mathrm{TOL}$ denotes the prescribed target tolerance. The smoothing parameters $c$ and $d$ should ideally be selected as functions of $\mathrm{TOL}$ to preserve the unbiasedness of the estimator for $q_{w}$. Although developing such an adaptive smoothing strategy is ongoing work, this paper fixes the smoothing parameters to $c=0.5$ and $d=0.125$. This simplification allows the analysis of the computational work of the MLIS method as a function of $\mathrm{TOL}$ without introducing additional complexity from the dependence of $\Delta^{c,d}$ on $\mathrm{TOL}$.
@@
The MLIS estimator (48) employs different likelihoods for levels $\ell$ and $\ell-1$. Therefore, the variance of level differences in MLIS takes the following form:
$V^{\mathrm{IS}}_{\ell,\ell-1}=\operatorname{Var}\left[g_{w}\left(\tilde{Z}^{\ell}_{N_{\ell}}\right)L_{N_{\ell}}-g_{w}\left(\tilde{Z}^{\ell-1}_{N_{\ell-1}}\right)L_{N_{\ell-1}}\right].$ (58)
@@
Figure 6: Smooth case with single-level likelihood. Left: Variance $V^{\mathrm{IS}}_{\ell+1,\ell}$. Right: Ratio $\frac{V^{\mathrm{IS}}_{\ell+1,\ell}}{V_{{\ell}}^{\mathrm{IS}}}$.
@@ 5.2 Common Likelihood
Next, this work proposes a novel variant of the MLIS framework that slightly departs from the standard approach. Instead of computing the optimal control independently at levels $\ell$ and $\ell-1$, this approach computes the control only at level $\ell$ and reuses it at level $\ell-1$. In other words, the controlled process $\tilde{Z}_{N_{\ell}}^{\ell}$ follows the same dynamics as those previously described in (51), and the coarse path $\tilde{Z}_{N_{\ell-1}}^{\ell-1}$ is defined using a coupled construction.
The shifted Brownian increment at level $\ell$ is denoted by $\Delta\tilde{W}_{n}^{\ell}=\Delta W_{n}^{\ell}+\zeta^{*}(t_{n}^{\ell},\tilde{\boldsymbol{X}}_{n}^{\ell},\tilde{Z}_{n}^{\ell})\,\Delta t_{\ell}$ and is a Brownian increment under a new measure by Girsanov’s theorem [30]. Then, the coarse-level increment is defined as the sum of two consecutive fine-level increments:
@@
Unlike the single-level likelihood approach, where the fine and coarse paths are coupled prior to the change of measure, the present method performs the coupling after the change of measure. This strategy ensures that the same likelihood $L_{N_{\ell}}$ is employed for levels $\ell$ and $\ell-1$. Under this construction, the variance of the multilevel difference becomes the following:
$\tilde{V}^{\mathrm{IS}}_{\ell,\ell-1}=\operatorname{Var}\left[\left(g_{w}\left(\tilde{Z}_{N_{\ell}}^{\ell}\right)-g_{w}\left(\tilde{Z}_{N_{\ell-1}}^{\ell-1}\right)\right)L_{N_{\ell}}\right].$ (62)
In the smooth case, this variance is expected to decay at the same rate (rate = 2) as the standard MLMC variance. Figure 7 plots $\tilde{V}^{\mathrm{IS}}_{\ell+1,\ell}$ along with the ratio $\frac{\tilde{V}^{\mathrm{IS}}_{\ell+1,\ell}}{V_{{\ell}}^{\mathrm{IS}}}$. When using a common likelihood, the decay rate of $\tilde{V}^{\mathrm{IS}}_{\ell+1,\ell}$ improves to an order of 2, and the necessary condition (46) is satisfied for smaller values of $\ell_{0}$ compared with the single-level likelihood case. This result suggests that MLIS can outperform SLIS for reasonably small tolerances.
Figure 7: Smooth case using the common likelihood. Left: Variance $\tilde{V}^{\mathrm{IS}}_{\ell+1,\ell}$. Right: Ratio $\frac{\tilde{V}^{\mathrm{IS}}_{\ell+1,\ell}}{V_{{\ell}}^{\mathrm{IS}}}$.
The condition (46) is necessary but not sufficient; hence, this work proceeds to verify the condition in (35) to determine whether the improved variance decay rate is sufficient for MLIS to outperform SLIS. Figure 8 plots both sides of the inequality (35) to identify the values of $\ell_{0}$ for which the condition is satisfied.
Figure 8: Left-hand side $\mathrm{(LHS)}=\sqrt{\frac{V_{\ell_{0}+1,\ell_{0}}}{V_{{\ell_{0}}}}}$ and right-hand side $\mathrm{(RHS)}=\frac{\sqrt{2\frac{V_{{\ell_{0}+1}}}{V_{{\ell_{0}}}}}-1}{\sqrt{3}}$ by auxiliary HJB-PDE accuracy.
The sufficient condition is satisfied for the following:
@@
These results were numerically verified by plotting the sampling work of SLIS and MLIS for various auxiliary HJB-PDE accuracy values in Figure 9. As expected, MLIS requires less sampling work than SLIS for tolerances below the corresponding threshold $\mathrm{TOL}_{\text{lim}}^{(P)}$ for each auxiliary HJB-PDE resolution. These results confirm Proposition 1: the necessary and sufficient condition in (35) accurately predicts when MLIS outperforms SLIS. For each auxiliary HJB-PDE accuracy value, MLIS becomes more efficient once the tolerance drops below $\mathrm{TOL}_{\text{lim}}^{(P)}$.

### v2

@@ Abstract
This study considers the estimation of the complementary cumulative distribution function of the occupation time (i.e., the time spent below a threshold) for a process governed by a stochastic differential equation. The focus is on the right tail, where the underlying event becomes rare, and using variance reduction techniques is essential to obtain computationally efficient estimates. Building on recent developments that relate importance sampling (IS) to stochastic optimal control, this work develops an optimal single-level IS (SLIS) estimator based on the solution of an auxiliary Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE). The cost of solving the HJB-PDE is incorporated into the total computational work, and an optimized trade-off between preprocessing and sampling is proposed to minimize the overall cost. The SLIS approach is extended to the multilevel setting to enhance efficiency, yielding a multilevel IS (MLIS) estimator. A key feature that emerges in the MLIS framework is that the single-level variance decreases to zero as the discretization level increases, a behavior that stems from the zero-variance property of the optimal control. This observation leads to the formulation of a necessary and sufficient condition under which MLIS outperforms SLIS. To ensure that this condition is satisfied, we introduce a smoothing of both the drift and the observable to enhance the variance decay rate, together with a novel common-likelihood MLIS formulation that preserves this decay under IS. The classical multilevel Monte Carlo complexity theory can be extended to accommodate settings in which the variance on coarse levels depends on the target accuracy, a phenomenon for which the variance-decay behavior in the IS setting is a key example. Notably, the total work complexity of MLIS can be better than than quadratic. Numerical experiments in the context of fade duration estimation demonstrate the benefits of the proposed approach and validate the theoretical results, including the analysis of the smoothing error, which is shown to be negligible. Consequently, the smoothed MLIS estimator remains unbiased for the occupation time while achieving significantly better performance than SLIS.
Keywords: Rare-event estimation, occupation time, stochastic differential equation, importance sampling, multilevel Monte Carlo, Hamilton–Jacobi–Bellman, variance decay, smoothing error.
@@ 1 Introduction
This study considers a system in which a process evolves according to a stochastic differential equation (SDE), with the goal of characterizing the distribution of the occupation time, specifically, the time a process spends below a critical threshold. Occupation time has been analytically studied in the context of Markov chains using matrix-based techniques [31]. In contrast, the present work addresses a more general setting in which the underlying dynamics are governed by SDEs with continuous state spaces, requiring advanced stochastic-simulation techniques for accurate estimation.
@@
Among such techniques, importance sampling (IS) [25, 12] is widely used for rare-event estimation in SDEs. Several studies have established a connection between IS and stochastic optimal control (SOC) [22, 6, 33, 10, 1, 21, 30], where the optimal change of measure is characterized by the solution of an auxiliary Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE). For example, [10] develops an optimal change of measure via SOC to estimate rare-event probabilities for McKean–Vlasov SDEs, while [1] applies a similar approach to estimate the CCDF of the fade duration in wireless communications. Although these studies have demonstrated substantial variance reduction and have analyzed the sampling cost of IS estimators, they do not account for the computational cost of obtaining the optimal change of measure because they solve the required auxiliary HJB-PDE off-line.
In the present study, the trade-off between the accuracy of the control (12) and the resulting variance reduction is investigated, and the cost of solving the auxiliary HJB-PDE (13) is incorporated into a comprehensive analysis of the total computational work of the IS estimator. The analysis reveals that the variance of the IS estimator decreases with the target accuracy, resulting in a lower overall computational cost for IS compared to the crude MC method.
In order to further improve the efficiency of the single-level IS (SLIS) estimator for rare-event probability estimation, we extend our previous work [1] to the multilevel setting by combining IS with the multilevel MC (MLMC) method introduced by [19]. Several studies have investigated combining IS with the MLMC method. For example, one study [8] applied IS to address the high‑kurtosis phenomenon in the MLMC method for stochastic reaction networks, and another study [5, 23] proposed variance‑reduction schemes based on suboptimal, levelwise constant controls. The most recent development [11] combines SOC-based IS with the MLMC method in a single‑likelihood multilevel IS (MLIS) framework, achieving substantial computational savings. In the context of Bayesian inverse problems, various methods, such as those in [17, 28], exploit coarse-level likelihood evaluations to inform fine-level sampling, demonstrating that reusing the likelihood structure across model hierarchies can be highly effective. Building on these insights and following the method in [11], this work introduces a novel common-likelihood MLIS approach. This formulation preserves the variance-convergence rate of MLMC and reduces the total computational work required by MLIS.
@@
A further contribution of this work is the formulation of a necessary and sufficient condition under which MLIS outperforms SLIS. To ensure that this condition is satisfied in the occupation time case, we introduce a smoothing of both the drift (3) and the observable (5), which enhances the decay rate of the multilevel variance. Moreover, we show that the proposed common-likelihood MLIS formulation preserves this improved decay under IS. Since smoothing modifies the original problem, we also analyze the induced smoothing error and prove that it is negligible, ensuring that the resulting smoothed MLIS estimator remains unbiased for the occupation time.
@@
  • In the setting of occupation-time estimation, a smoothing procedure for both the drift and the observable is introduced to improve the variance decay and ensure that the necessary condition for MLIS to outperform SLIS is satisfied.
  • The resulting smoothing error is analyzed and shown to be negligible, guaranteeing that the proposed smoothed MLIS estimator remains unbiased while achieving a lower computational cost than SLIS.
The remainder of this paper is structured as follows. Section 2 describes the problem setting. Next, Section 3 details the SLIS method, presenting the construction of the optimal estimator, exploring the influence of the accuracy of the auxiliary HJB-PDE on the variance, and reporting the numerical results. Then, Section 4 reviews the MLMC method and provides the sufficient and necessary conditions for its efficiency. Section 5 introduces the MLIS approach, including the single-level likelihood formulation, the smoothing strategy, and the common-likelihood construction. Moreover, Section 6 presents the analysis of the effect of the HJB-PDE accuracy on the efficiency of the MLIS method and the proposed optimal work formulation, providing a numerical example and investigating the work rate when the single-level variance decays. Finally, Section 7 analyzes the impact of smoothing on the estimator, derives a quadratic characterization of the smoothing error, and establishes parameter-selection strategies ensuring that smoothing introduces only a negligible bias in the occupation-time setting.
@@ 2 Problem Setting
where
$f(x):=\mathbbm{1}_{\{x<\gamma_{\text{th}}\}}.$ (3)
The aim is to characterize the CCDF of the occupation time $Z(T)$, that is,
@@
where
$g_{w}(x)=\mathbbm{1}_{\{x>w\}}\qquad\text{for a given threshold $w>0$.}$ (5)
This formulation can be compactly described by the following coupled system:
@@ 4.2 Sufficient and Necessary Condition for MLMC Advantages
See Appendix A. ∎
@@
See Appendix B. ∎
@@ 5.1 Single-Level Likelihood
$\mathcal{A}_{\mathrm{MLIS}}=\frac{1}{M_{0}}\sum_{m=1}^{M_{0}}\tilde{g}_{w}^{(0,m)}+\sum_{\ell=1}^{L}\frac{1}{M_{\ell}}\sum_{m=1}^{M_{\ell}}\left(\tilde{g}_{w}^{\ell}-\tilde{g}_{w}^{\ell-1}\right)^{(m)},\qquad\text{where}\quad\tilde{g}_{w}^{\ell}=g_{w}\left(\tilde{Z}_{N_{\ell}}^{\ell}\right)L_{N_{\ell}}.$ (41)
@@ 5.1.1 Nonsmooth Case
The same application described in Section 3.3 is considered, where the observable $g_{w}$ and drift $f$ are nonsmooth functions. This work investigates the performance of the proposed MLIS estimator (41) by fixing an arbitrary coarse level $N_{0}=20$ and computing the variance $V^{\mathrm{IS}}_{\ell,\ell-1}$ for values of $\ell$ to assess whether the condition in (39) can be satisfied. This analysis is repeated for various levels of the auxiliary HJB-PDE accuracy, following the same approach as earlier to study the behavior of $V_{N}^{\mathrm{IS}}$ (see Section 3.3).
@@
Figure 4 presents the results and reveals that no prospect exists for MLIS to outperform SLIS when $P\geq 80$. The only scenario in which MLIS may offer a slight advantage is when $P=40$. Even then, the advantage is only for tolerances requiring numerous levels, specifically, when $\mathrm{TOL}\leq 0.003$, corresponding to $L_{\mathrm{opt}}\geq 8$. We compute the sampling work of MLIS using the arbitrary choice $N_{0}=20$ to further investigate this observation. Over the range of relative tolerances $[0.1,0.003]$, SLIS consistently outperforms MLIS for all auxiliary HJB-PDE accuracy values. For a more rigorous comparison, the optimal coarse level $\ell_{0}^{\mathrm{opt}}$ was computed using Algorithm 2, demonstrating that $\ell_{0}^{\mathrm{opt}}=L_{\mathrm{opt}}$ for all considered tolerance values, implying that MLIS reduces to SLIS across the entire tolerance range.
@@ 5.1.2 Smooth Case
This section introduces the smooth approximations $f^{d}$ and $g_{w}^{c}$ for the functions $f$ and $g_{w}$, respectively, to improve the multilevel variance convergence rates. Rather than estimating the original quantity $q_{w}$, the smoothed quantity $q_{w}^{c,d}=\mathbb{E}\left[g_{w}^{c}(Z^{d}(T))\right]$ is approximated, where $Z^{d}(T)=\int_{0}^{T}f^{d}\left(h(\boldsymbol{X}(s))\right)\,ds$.
Ideally, the smoothing parameters $c$ and $d$ should be chosen as a function of $\mathrm{TOL}$ in order to retain the effective unbiasedness of the estimator for $q_{w}$. The construction of such a tolerance dependent strategy is developed in Section 7.2. In the following, however, we fix the smoothing parameters to $c=0.5$ and $d=0.122$ for the entire range of $\mathrm{TOL}$. This choice will be justified in Section 7.4, where it is shown to yield a negligible smoothing error. Consequently, the analysis of the computational work of the MLIS method as a function of $\mathrm{TOL}$ in Figure 9 and 10 is performed without the additional complication of a tolerance dependent smoothing error.
@@
We now consider the smoothed dynamics obtained by replacing $f$ with $f^{d}$ in (43)–(45). We denote the corresponding auxiliary processes by $\tilde{Z}^{\ell,d}$ and $\tilde{Z}^{\ell-1,d}$. In the smoothed setting, the likelihood weights defined in (47)–(48) also depend on the parameters $(c,d)$, since both the control $\zeta^{*}$ and the process $\tilde{Z}$ are replaced by their smoothed counterparts. We therefore write them as $L_{N_{\ell}}^{c,d}$ and $L_{N_{\ell-1}}^{c,d}$ in what follows. The MLIS estimator (41) employs different likelihoods for levels $\ell$ and $\ell-1$. Therefore, the variance of level differences in MLIS in the smooth case takes the following form:
$V^{\mathrm{IS}}_{\ell,\ell-1}(c,d)=\operatorname{Var}\left[g_{w}^{c}\left(\tilde{Z}^{\ell,d}_{N_{\ell}}\right)L_{N_{\ell}}^{c,d}-g_{w}^{c}\left(\tilde{Z}^{\ell-1,d}_{N_{\ell-1}}\right)L_{N_{\ell-1}}^{c,d}\right].$ (50)
@@
Figure 6: Smooth case ($d=0.122$ and $c=0.5$) with single-level likelihood. Left: Variance $V^{\mathrm{IS}}_{\ell+1,\ell}$. Right: Ratio $\frac{V^{\mathrm{IS}}_{\ell+1,\ell}}{V_{{\ell}}^{\mathrm{IS}}}$.
@@ 5.2 Common Likelihood
Next, this work proposes a novel variant of the MLIS framework that slightly departs from the standard approach. Instead of computing the optimal control independently at levels $\ell$ and $\ell-1$, this approach computes the control only at level $\ell$ and reuses it at level $\ell-1$. In other words, the controlled process $\tilde{Z}_{N_{\ell}}^{\ell}$ follows the same dynamics as those previously described in (43), and the coarse path $\tilde{Z}_{N_{\ell-1}}^{\ell-1}$ is defined using a coupled construction. The shifted Brownian increment at level $\ell$ is denoted by $\Delta\tilde{W}_{n}^{\ell}=\Delta W_{n}^{\ell}+\zeta^{*}(t_{n}^{\ell},\tilde{\boldsymbol{X}}_{n}^{\ell},\tilde{Z}_{n}^{\ell})\,\Delta t_{\ell}$ and is a Brownian increment under a new measure by Girsanov’s theorem [29]. Then, the coarse-level increment is defined as the sum of two consecutive fine-level increments:
@@
Unlike the single-level likelihood approach, where the fine and coarse paths are coupled prior to the change of measure, the present method performs the coupling after the change of measure. This strategy ensures that the same likelihood $L_{N_{\ell}}$ is employed for levels $\ell$ and $\ell-1$. Under this construction, the variance of the multilevel difference in the smooth case becomes the following:
$\tilde{V}^{\mathrm{IS}}_{\ell,\ell-1}(c,d)=\operatorname{Var}\left[\left(g_{w}^{c}\left(\tilde{Z}_{N_{\ell}}^{\ell,d}\right)-g_{w}^{c}\left(\tilde{Z}_{N_{\ell-1}}^{\ell-1,d}\right)\right)L_{N_{\ell}}^{c,d}\right].$ (54)
This variance is expected to decay at the same rate (rate = 2) as the standard MLMC variance.
Figure 7: Smooth case ($d=0.122$ and $c=0.5$) using the common likelihood. Left: Variance $\tilde{V}^{\mathrm{IS}}_{\ell+1,\ell}$. Right: Ratio $\frac{\tilde{V}^{\mathrm{IS}}_{\ell+1,\ell}}{V_{{\ell}}^{\mathrm{IS}}}$.
Figure 8: Left-hand side $\mathrm{(LHS)}=\sqrt{\frac{\tilde{V}^{\mathrm{IS}}_{\ell_{0}+1,\ell_{0}}}{\tilde{V}^{\mathrm{IS}}_{{\ell_{0}}}}}$ and right-hand side $\mathrm{(RHS)}=\frac{\sqrt{2\frac{\tilde{V}^{\mathrm{IS}}_{{\ell_{0}+1}}}{\tilde{V}^{\mathrm{IS}}_{{\ell_{0}}}}}-1}{\sqrt{3}}$.
Figure 7 plots $\tilde{V}^{\mathrm{IS}}_{\ell+1,\ell}$ along with the ratio $\frac{\tilde{V}^{\mathrm{IS}}_{\ell+1,\ell}}{V_{{\ell}}^{\mathrm{IS}}}$. When using a common likelihood, the decay rate of $\tilde{V}^{\mathrm{IS}}_{\ell+1,\ell}$ improves to an order of 2, and the necessary condition (39) is satisfied for smaller values of $\ell_{0}$ compared with the single-level likelihood case. This result suggests that MLIS can outperform SLIS for reasonably small tolerances.
The condition (39) is necessary but not sufficient; hence, this work proceeds to verify the condition in (37) to determine whether the improved variance decay rate is sufficient for MLIS to outperform SLIS. Figure 8 plots both sides of the inequality (37) to identify the values of $\ell_{0}$ for which the condition is satisfied. The sufficient condition is satisfied for the following:
@@
These results were numerically verified by plotting the sampling work of SLIS and MLIS in the smooth case for various auxiliary HJB-PDE accuracy values in Figure 9. As expected, MLIS requires less sampling work than SLIS for tolerances below the corresponding threshold $\mathrm{TOL}_{\text{lim}}^{(P)}$ for each auxiliary HJB-PDE resolution. These results confirm Proposition 1: the necessary and sufficient condition in (37) accurately predicts when MLIS outperforms SLIS. For each auxiliary HJB-PDE accuracy value, MLIS becomes more efficient once the tolerance drops below $\mathrm{TOL}_{\text{lim}}^{(P)}$.